# When Data is the Algorithm: A Systematic Study and Curation of Preference Optimization Datasets

### v1

@@ 1 Introduction
Learning from preference feedback or commonly called Reinforcement Learning from Human Feedback (RLHF) is an important final step in aligning large language models (LLMs). This stage refines LLMs to enhance their performance on a wide range of downstream capabilities including instruction following, math, and code (wang2023aligning; shen2023large).
Within preference tuning, two prominent approaches are reward-based and reward-free. Reward-based methods train a reward model from preference data and optimize it using algorithms such as Proximal Policy Optimization (PPO) (schulman2017proximalpolicyoptimizationalgorithms). In contrast, reward-free methods, including Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage), remove the need for an explicit reward function. DPO, in particular, has become a favored technique for its efficiency and simplicity: models are directly trained to prefer one completion over another using curated preference pairs, without requiring a reward model or policy rollouts (ivison2024unpackingdpoppodisentangling).
However, proprietary LLMs leverage vast undisclosed DPO corpora that remain inaccessible to practitioners, often because of licensing restrictions or intellectual property concerns. Nevertheless, the open-source community has made notable progress by releasing open DPO datasets, including TuluDPO (lambert2025tulu3pushingfrontiers), ORPO (orpo), UltraFeedback (cui2024ultrafeedbackboostinglanguagemodels), HelpSteer (helpsteer), and Code-Preference-Pairs (codepreferences). These corpora have become essential building blocks for many open models and contribute meaningfully to the broader reproducibility efforts in LLM preference tuning (lambert2025tulu3pushingfrontiers; bakouch2025smollm3; allal2025smollm2smolgoesbig).
@@
Another challenge is that while some datasets report their coarse compositions, they remain opaque at the sample level: Quality annotations are mostly missing and only a few datasets provide explicit human-annotated preference scores to justify chosen completions, while others contain only binary pairs without ranking information, leaving it unclear how much better the preferred sample actually is. This lack of clarity hampers progress by making it difficult to design systematic curation recipes.
@@
  • Comparative Evaluation: We present the first systematic cross-analysis of five open DPO datasets: TuluDPO, ORPO, UltraFeedback, HelpSteer, and Code-Preference-Pairs, spanning both general-purpose and domain-specific tasks. We conduct preference fine-tuning on eight different models of various scales and evaluate performance across 12 common benchmarks from the popular Open LLM Leaderboards (leaderboardv1; open-llm-leaderboard-v2) and two additional code generation benchmarks. By holding all training parameters constant, we allow for a fair comparison between DPO datasets.
  • Sample-Level Annotations and Analysis: Using the Magpie framework (xu2024magpie), we annotate each preference pair with metadata for task category, difficulty, input quality, and preference reward, an independent, reward-model-based signal that helps evaluate whether chosen completions are indeed better justified than discarded ones. This allows us to assess preference signals at scale and to validate the reliability of the preference order, particularly when human annotations are unavailable or inconsistent. In our extensive analysis, we reveal substantial variation in dataset composition, quality, and alignment performance, both across tasks and model scales.
@@
Our work provides practical guidance for curating effective data mixtures from open-source corpora. To support ongoing research in preference optimization, we publicly release our annotated versions of TuluDPO [^1] Annotated TuluDPO dataset: huggingface.co/datasets/aladinDJ/tulu-DPO-annotated, ORPO [^2] Annotated ORPO dataset: huggingface.co/datasets/aladinDJ/orpo-DPO-annotated, UltraFeedback [^3] Annotated UltraFeedback dataset: huggingface.co/datasets/aladinDJ/ultrafeedback-DPO-annotated, HelpSteer [^4] Annotated HelpSteer dataset: huggingface.co/datasets/aladinDJ/helpsteer-DPO-annotated, Code-Preference-Pairs [^5] Annotated Codepreferences dataset: huggingface.co/datasets/aladinDJ/codepreferences-DPO-annotated, as well as UltraMix [^6] Annotated UltraMix dataset: huggingface.co/datasets/aladinDJ/ultramix-DPO-annotated.
@@ 2 Background and Motivation
DPO is an offline RL approach that learns directly from preference feedback by optimizing policies on curated preference pairs. As a result, the quality and diversity of preference data are critical for improving model performance across a wide range of capabilities. Although several DPO mixtures have been released in the literature, no systematic performance comparisons exist. In particular, prior work lacks comprehensive, side-by-side evaluations of sample quality and task coverage, even though such a comparison is critical for designing principled and effective DPO mixtures.
We close these gaps by presenting the first systematic analysis of five widely used open DPO datasets: two general-purpose (TuluDPO, ORPO), two instruction-following (UltraFeedback, HelpSteer), and one code-focused (Code-Preference-Pairs). See App. B for details on these datasets. Our efforts are inspired by work (tulutalk), which systematically annotated and analyzed SFT datasets.
While several preference datasets exist, we selected these five for their frequent appearance in research papers and open-source dataset hubs (mlabonne_llm-datasets_2025). However, our analysis, annotation pipeline, and data curation recipes are generalizable and can, in principle, be applied to any DPO dataset. In the following sections, we benchmark the performance of each corresponding DPO dataset and conduct an extensive side-by-side comparison of their compositions on a per-sample basis.
@@ 3 Benchmarking DPO Datasets
We begin with a side-by-side performance comparison of the five DPO datasets in Table 1. We perform DPO training using Open-Instruct (allenai_openinstruct) on two representative models: (i) Llama-3.1-8B-TuluSFT, an SFT-tuned Llama-3.1-8B-Base (grattafiori2024llama3), and (ii) Qwen-2.5-7B-TuluSFT, an SFT-tuned Qwen-2.5-7B-Base (yang2024qwen2). Here SFT is performed using the popular TuluSFT dataset (lambert2025tulu3pushingfrontiers). We benchmark downstream performance on 12 representative tasks from Open LLM Leaderboards V1 and V2 (leaderboardv1; open-llm-leaderboard-v2), as well as on two code generation tasks: HumanEval and HumanEval+ (chen2021codex). We use this setup for all of our experiments. More details on training and evaluation are found in App. F.1. We further extend our evaluations to six additional open models in Sec. 5.3 with comprehensive results in App. F.3.
@@ 4.1 Annotating DPO Samples using Magpie
Magpie leverages an LLM as a judge to assign labels for task category (12 classes), query difficulty (rated from very easy to very hard), input quality (rated from very poor to excellent), safety, and language, using specialized prompt templates (see App. C.1). As a novel component, we assess preference rewards by evaluating the response quality of chosen and discarded completions using FsfairX, a Llama-3-8B-Instruct-based reward model introduced by reward_model, which was fine-tuned on a diverse mixture of high-quality preference datasets. This particular reward model has proven to achieve strong alignment with human preference judgments, making it a reliable signal for quantifying how well the model completions satisfy the instruction. To avoid over-reliance on a single reward model, we include a brief comparison with an alternative model in App. C.2, yielding similar principal outcomes and thereby validating our choice. For all other labels, we use Llama-3.3-70B-Instruct as the judge model given its demonstrated reliability, particularly for difficulty and quality annotations, as assessed by tulutalk. In the following analysis, we focus on task category, query difficulty, input quality, and preference reward as the primary signals for dataset comparison, and refer to App. D.5 for additional details on low-impact language and safety distributions.
@@ 4.3 Preference Reward Analysis
As we compute preference rewards with an independent reward model, it also serves as a verification mechanism for the overall preference ordering, which is particularly interesting in the case of human- or GPT-based annotations as found in UltraFeedback and HelpSteer. To this end, we use the preference reward annotations to first evaluate whether the chosen completion in each preference pair indeed receives a higher reward than the rejected one. Surprisingly, we observe a non-negligible number of cases where the reward model assigns a higher score to the discarded response. Fig. 5 shows, for each DPO dataset, the proportion of samples where the chosen completion is correctly preferred according to the reward model. TuluDPO, ORPO, UltraFeedback, and HelpSteer exhibit similar patterns, with only 70–80% of samples aligning with the chosen preference. This suggests that preference decisions may sometimes be arbitrary or based on near-identical completions (see App. B). In addition, this shows that GPT-4-based reward annotations in UltraFeedback, as well as the simplistic use of mean scores in HelpSteer, are often misaligned with the judgments of a specialized preference reward model. Interestingly, Code-Preference-Pairs appears less affected by such inconsistencies. This may be due to more salient signals (e.g., missing features or syntax errors) in code tasks, making it easier for both humans and reward models to distinguish between preference pairs.
@@
From this analysis, we conclude that our preference reward annotations uncover several non-trivial insights into dataset quality and curation practices. Most notably, our independent reward model is able to identify potentially misaligned preference pairs. Further, our findings highlight that both input quality as well as consistency between chosen and discarded completions emerge as useful filtering criteria, motivating the construction of potentially more performant DPO mixtures by selectively retaining samples with high-quality prompts and coherent preference ordering. We refer to App. D.4 for additional results and insights.
@@ 5.1 Quality- and Reward-Based Curation Recipe
Task Analysis. Our initial mixture underperforms on benchmarks that heavily rely on instruction following capabilities. This is a critical skill that has been shown to influence performance across a wide range of tasks, as observed by lambert2025tulu3pushingfrontiers, tulutalk, and cohere2025commanda. Since the majority of samples in our mixtures originate from TuluDPO (due to its significantly larger size compared to the other DPO datasets), we examine the distribution of instruction following task categories in Table 3. The analysis reveals that UltraMix-170k underrepresents information seeking and reasoning samples, with relative reductions of 20% and 13%, respectively. In contrast, the proportions of math and coding samples have increased. However, as previously noted, performance on these categories tends to depend on absolute sample count than relative proportion. These findings confirm that our quality- and reward-based curation recipe would benefit from task-aware filtering to preserve a stronger representation of instruction following tasks that increase overall performance.
@@ 5.2 Quality-, Reward-, and Task-Based Curation Recipe
Compared to UltraMix-170k, we observe substantial improvements for both new mixtures on IFEval, GSM8K, MATH, and coding, suggesting that the inclusion of more samples has increased overall performance across benchmarks. While UltraMix-187k still underperforms TuluDPO on IFEval and MATH for Llama, the additional 3,000 instruction following samples in UltraMix-190k have led to a compounding performance increase across all benchmarks. This corroborates the observations by lambert2025tulu3pushingfrontiers, tulutalk, and cohere2025commanda that instruction following samples can lead to noticeable cross-task performance improvements.
@@
Collectively, these results show that UltraMix-190k consistently achieves top-tier performance across a diverse set of benchmarks, offering substantial efficiency gains with 30% fewer samples than TuluDPO. We designate UltraMix-190k as our final UltraMix DPO dataset and refer to App. F.5 for a dedicated analysis on training efficiency. To assess the role of reward filtering, we evaluate UltraMix curated without it in App. F.4, showing that quality- and task-based filters alone are insufficient and confirming the need and effectiveness of our principled reward-based data curation recipe.
@@ 5.3 Generalization to Different Architectures and Model Scales
To demonstrate the effectiveness of UltraMix, we evaluate six additional open-source models of different architectures and scales: large models (Apertus-8B-SFT (swissai2025apertus), OLMo-2-7B-SFT (olmo20242olmo2furious)), medium models (SmolLM-3-3B-SFT (bakouch2025smollm3), Instella-3B-SFT (Instella)), and small models (SmolLM-2-1.7B-SFT (allal2025smollm2smolgoesbig), OLMo-2-1B-SFT (olmo20242olmo2furious)). In contrast to our earlier experiments with TuluSFT-tuned Llama and Qwen variants, these additional models have publicly released their SFT checkpoints, which enables an important axis for evaluation with instruct models that were SFT-tuned using different datasets.
@@ 6 Conclusion
In this work, we systematically analyzed five widely used DPO post-training datasets: TuluDPO, ORPO, UltraFeedback, HelpSteer, and Code-Preference-Pairs. By annotating each preference pair with labels for input quality, difficulty, and preference reward, we identified high-reward, high-quality samples and developed a principled data curation strategy grounded in insights from targeted ablation studies. Our curation strategy led to the construction of UltraMix, a new DPO mixture that is 30% smaller than TuluDPO while achieving superior performance across a broad set of evaluation benchmarks. The results presented in this study support several key takeaways: (a) judicious incorporation of input quality, difficulty, task category, and preference rewards leads to systematic curation recipes, (b) coherent reward samples play a critical role in driving performance gains in DPO training, (c) the optimal composition of data mixtures is task-sensitive, requiring careful trade-offs between quality and task diversity, particularly for instruction-following tasks. We validate the competitiveness of UltraMix by conducting extensive evaluations across diverse benchmarks and eight LLMs of different sizes. By illustrating how quality-, reward-, and task-aware curation can reduce data requirements while increasing performance, our work provides a strong foundation for future efforts in systematic, reward-aligned DPO dataset design. We discuss limitations and broader impact in App. G.

### v2

# arXiv:2511.10985v2[cs.CL] 01 Mar 2026

@@ 1 Introduction
Learning from preference feedback or commonly called Reinforcement Learning from Human Feedback (RLHF) is an important final step in aligning large language models (LLMs). This stage refines LLMs to enhance their performance on a wide range of downstream capabilities, including instruction following, math, and code (Wang et al., 2023; Shen et al., 2023).
Within preference tuning, two prominent approaches are reward-based and reward-free alignment. Reward-based methods train a reward model from preference data and optimize it using algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017). In contrast, reward-free methods, including Direct Preference Optimization (DPO) (Rafailov et al., 2023), remove the need for an explicit reward function. DPO, in particular, has become a favored technique for its efficiency and simplicity: models are directly trained to prefer one completion over another using curated preference pairs, and thus no reward model or policy rollouts are required (Ivison et al., 2024).
However, proprietary LLMs leverage vast undisclosed DPO corpora that remain inaccessible to practitioners, often because of licensing restrictions or intellectual property concerns. Nevertheless, the open-source community has made notable progress by releasing open DPO datasets, including TuluDPO (Lambert et al., 2025), ORPO (Labonne, 2024), UltraFeedback (Cui et al., 2024), HelpSteer (Wang et al., 2024c), and Code-Preference-Pairs (Vezora, 2024). These corpora have become essential building blocks for many open models and contribute meaningfully to the broader reproducibility efforts in LLM preference tuning (Lambert et al., 2025; Bakouch et al., 2025; Allal et al., 2025).
@@
Another challenge is that while some datasets report their coarse compositions, they remain opaque at the sample level: rich annotations are mostly missing and only a few datasets provide explicit human-annotated preference scores to justify chosen completions, while others contain only binary pairs without ranking information, leaving it unclear how much better the preferred sample actually is. This lack of clarity hampers progress by making it difficult to design systematic curation recipes.
@@
  • Comparative Evaluation: We present the first systematic cross-analysis of five open DPO datasets: TuluDPO, ORPO, UltraFeedback, HelpSteer, and Code-Preference-Pairs, spanning both general-purpose and domain-specific tasks. We conduct preference fine-tuning on eight different models of various scales and evaluate performance across 12 common benchmarks from the popular Open LLM Leaderboards (Fourrier et al., 2023; 2024) and two additional code generation benchmarks. By holding all training parameters constant, we allow for a fair comparison between DPO datasets.
  • Sample-Level Annotations and Analysis: Using the Magpie framework (Xu et al., 2025), we annotate each preference pair with metadata for task category, difficulty, input quality, and preference reward, an independent, reward-model-based signal that helps evaluate whether chosen completions are indeed better justified than discarded ones. This allows us to assess preference signals at scale and to validate the reliability of the preference order, particularly when human annotations are unavailable or inconsistent. In our extensive analysis, we reveal substantial variation in dataset composition, quality, and alignment performance, both across tasks and model scales.
@@
Our extensive study shows several interesting findings: a) misaligned preference orders (i.e., the chosen answer is not always better) are common and degrade performance, b) prompt quality and preference rewards are strongly correlated, and c) filtering based on a single signal (prompt quality or preference reward) is insufficient and requires combining multiple signals for effective curation.
Our work provides practical guidance for curating effective data mixtures from open-source corpora. To support ongoing research in preference optimization, we publicly release our annotated versions of TuluDPO [^1] Annotated TuluDPO dataset: huggingface.co/datasets/aladinDJ/tulu-DPO-annotated, ORPO [^2] Annotated ORPO dataset: huggingface.co/datasets/aladinDJ/orpo-DPO-annotated, UltraFeedback [^3] Annotated UltraFeedback dataset: huggingface.co/datasets/aladinDJ/ultrafeedback-DPO-annotated, HelpSteer [^4] Annotated HelpSteer dataset: huggingface.co/datasets/aladinDJ/helpsteer-DPO-annotated, Code-Preference-Pairs [^5] Annotated Codepreferences dataset: huggingface.co/datasets/aladinDJ/codepreferences-DPO-annotated, and UltraMix [^6] Annotated UltraMix dataset: huggingface.co/datasets/aladinDJ/ultramix-DPO-annotated.
@@ 2 Background and Motivation
DPO is an offline RL approach that learns directly from preference feedback by optimizing policies on curated preference pairs. As a result, the quality and diversity of preference data are critical for improving model performance across a wide range of capabilities. Although several DPO mixtures have been released, no systematic performance comparisons exist that enable fair assessment, i.e., evaluations that fix training across datasets to remove cross-study heterogeneity. In particular, prior work lacks comprehensive, side-by-side evaluations of sample quality and task coverage, even though such a comparison is critical for designing principled and effective DPO mixtures.
We close these gaps by presenting the first systematic analysis of five widely used open DPO datasets: two general-purpose (TuluDPO, ORPO), two helpfulness-oriented (UltraFeedback, HelpSteer), and one code-focused (Code-Preference-Pairs). See App. B.1 for details on these datasets. Our efforts are inspired by our previous work (Djuhera et al., 2025), which systematically annotated and analyzed datasets for the SFT case.
While several preference datasets exist, we selected these five for their frequent appearance in research papers and open-source dataset hubs (Labonne, 2025). However, our analysis, annotation pipeline, and data curation recipes are generalizable and can, in principle, be applied to any DPO dataset. We provide a thorough comparison with related work in App. B.2. In the following sections, we benchmark the performance of each corresponding DPO dataset and conduct an extensive side-by-side comparison of their compositions on a per-sample basis.
@@ 3 Benchmarking DPO Datasets
We begin with a side-by-side performance comparison of the five DPO datasets in Table 1. We perform DPO training using Open-Instruct (AllenAI, 2023) on two representative models: (i) Llama-3.1-8B-TuluSFT, an SFT-tuned Llama-3.1-8B-Base (Grattafiori et al., 2024), and (ii) Qwen-2.5-7B-TuluSFT, an SFT-tuned Qwen-2.5-7B-Base (Yang et al., 2025). Here SFT is performed using the popular TuluSFT dataset (Lambert et al., 2025). We benchmark downstream performance on 12 representative tasks from Open LLM Leaderboards V1 and V2 (Fourrier et al., 2023; 2024), as well as on two code generation tasks: HumanEval and HumanEval+ (Chen et al., 2021). We use this setup for all of our experiments. More details on training and evaluation are found in App. F.1. We further extend our evaluations to six additional open models in Sec. 5.3 with comprehensive results in App. F.3.
@@ 4.1 Annotating DPO Samples using Magpie
Magpie leverages an LLM as a judge to assign labels for task category (12 classes), query difficulty (rated from very easy to very hard), input quality (rated from very poor to excellent), safety, and language, using specialized prompt templates (see App. C.1). As a novel component, we assess preference rewards by evaluating the response quality of chosen and discarded completions using FsfairX, a Llama-3-8B-Instruct-based reward model introduced by Dong et al. (2024), which was fine-tuned on a diverse mixture of high-quality preference datasets. This particular reward model has proven to achieve strong alignment with human preference judgments, making it a reliable signal for quantifying how well the model completions satisfy the instruction. To avoid over-reliance on a single reward model, we include a brief comparison with an alternative model in App. C.2, yielding similar principal outcomes and thereby validating our choice. For all other labels, we use Llama-3.3-70B-Instruct as the judge model given its demonstrated reliability, particularly for difficulty and quality annotations, as assessed by Djuhera et al. (2025). In the following analysis, we focus on task category, query difficulty, input quality, and preference reward as the primary signals for dataset comparison, and refer to App. D.5 for additional details on low-impact language and safety distributions.
@@ 4.3 Preference Reward Analysis
As we compute preference rewards with an independent reward model, it also serves as a verification mechanism for the overall preference ordering, which is particularly interesting in the case of human- or GPT-based annotations as found in UltraFeedback and HelpSteer. To this end, we use the preference reward annotations to first evaluate whether the chosen completion in each preference pair indeed receives a higher reward than the rejected one. Surprisingly, we observe a non-negligible number of cases where the reward model assigns a higher score to the discarded response. Fig. 5 shows, for each DPO dataset, the proportion of samples where the chosen completion is correctly preferred according to the reward model. TuluDPO, ORPO, UltraFeedback, and HelpSteer exhibit similar patterns, with only 70–80% of samples aligning with the chosen preference. This suggests that preference decisions may sometimes be arbitrary or based on near-identical completions (see App. B.1). In addition, this shows that GPT-4-based reward annotations in UltraFeedback, as well as the simplistic use of mean scores in HelpSteer, are often misaligned with the judgments of a specialized preference reward model. Interestingly, Code-Preference-Pairs appears less affected by such inconsistencies. This may be due to more salient signals (e.g., missing features or syntax errors) in code tasks, making it easier for both humans and reward models to distinguish between preference pairs.
@@
From this analysis, we conclude that our preference reward annotations uncover several non-trivial insights into dataset quality and curation practices. Most notably, our independent reward model is able to identify potentially misaligned preference pairs. Further, our findings highlight that both input quality as well as consistency between chosen and discarded completions emerge as useful filtering criteria, motivating the construction of potentially more performant DPO mixtures by selectively retaining samples with high-quality prompts and coherent preference ordering. We refer to App. D.4 for additional analyses and to App. F.5 for more ablations on preference filtering in DPO datasets.
@@ 5.1 Quality- and Reward-Based Curation Recipe
Task Analysis. Our initial mixture underperforms on benchmarks that heavily rely on instruction following capabilities. This is a critical skill that has been shown to influence performance across a wide range of tasks, as observed by Lambert et al. (2025), Djuhera et al. (2025), and Aakanksha et al. (2025). Since the majority of samples in our mixtures originate from TuluDPO (due to its significantly larger size compared to the other DPO datasets), we examine the distribution of instruction following task categories in Table 3. The analysis reveals that UltraMix-170k underrepresents information seeking and reasoning samples, with relative reductions of 20% and 13%, respectively. In contrast, the proportions of math and coding samples have increased. However, as previously noted, performance on these categories tends to depend on absolute sample count than relative proportion. These findings confirm that our quality- and reward-based curation recipe would benefit from task-aware filtering to preserve a stronger representation of instruction following tasks that increase overall performance.
@@ 5.2 Quality-, Reward-, and Task-Based Curation Recipe
Compared to UltraMix-170k, we observe substantial improvements for both new mixtures on IFEval, GSM8K, MATH, and coding, suggesting that the inclusion of more samples has increased overall performance across benchmarks. While UltraMix-187k still underperforms TuluDPO on IFEval and MATH for Llama, the additional 3,000 instruction following samples in UltraMix-190k have led to a compounding performance increase across all benchmarks. This corroborates the observations by Lambert et al. (2025), Djuhera et al. (2025), and Aakanksha et al. (2025) that instruction following samples can lead to noticeable cross-task performance improvements.
@@
Collectively, these results show that UltraMix-190k consistently achieves top-tier performance across a diverse set of benchmarks, offering substantial efficiency gains with 30% fewer samples than TuluDPO. We designate UltraMix-190k as our final UltraMix DPO dataset and refer to App. F.6 for a dedicated analysis on training efficiency. To assess the contribution of individual filtering steps, we provide granular ablations in App. F.4, showing that quality-, task, and reward-based filters alone are insufficient and that UltraMix’s effectiveness stems from the principled combination of those signals.
@@ 5.3 Generalization to Different Architectures and Model Scales
To demonstrate the effectiveness of UltraMix, we evaluate six additional open-source models of different architectures and scales: large models (Apertus-8B-SFT (Apertus-Team, 2025), OLMo-2-7B-SFT (OLMo et al., 2024)), medium models (SmolLM-3-3B-SFT (Bakouch et al., 2025), Instella-3B-SFT (Liu et al., 2025)), and small models (SmolLM-2-1.7B-SFT (Allal et al., 2025), OLMo-2-1B-SFT (OLMo et al., 2024)). In contrast to our earlier experiments with TuluSFT-tuned Llama and Qwen variants, these additional models have publicly released their SFT checkpoints, which enables an important axis for evaluation with instruct models that were SFT-tuned using different datasets.
@@ 6 Conclusion
In this work, we systematically analyzed five widely used DPO post-training datasets: TuluDPO, ORPO, UltraFeedback, HelpSteer, and Code-Preference-Pairs. By annotating each preference pair with labels for input quality, difficulty, and preference reward, we identified high-reward, high-quality samples and developed a principled data curation strategy that led to UltraMix, a new DPO mixture that is 30% smaller than TuluDPO while achieving superior performance across benchmarks. The results presented in this study support several key takeaways: (a) judicious incorporation of input quality, difficulty, task category, and preference rewards leads to systematic curation recipes, (b) coherent reward samples play a critical role in driving performance gains in DPO training, and (c) the optimal composition of data mixtures is task-sensitive, requiring careful trade-offs between quality and task diversity. We validate the competitiveness of UltraMix by conducting extensive evaluations across 14 diverse benchmarks and eight LLMs of different sizes. By illustrating how quality-, reward-, and task-aware curation can reduce data requirements while increasing performance, our work provides a strong foundation for future efforts in systematic, reward-aligned DPO dataset design. We discuss limitations and broader impact in App. G.
Acknowledgments
This work was supported in part by the German Federal Ministry of Research, Technology and Space (BMFTR) within the research hub 6G-life (Grant 16KISK002), by the Bavarian Ministry of Science and the Arts and the Saxon Ministry for Science, Culture, and Tourism through the project Next Generation AI Computing (gAIn), by the Bavarian Ministry of Economic Affairs, Regional Development and Energy through the project 6G Future Lab Bavaria, and in part by IBM Research.